# Symmetry operations and Critical Behaviour in Classical to Quantum Stochastic Processes

### v1

@@ IV Numerical simulations
Previously we define several quantum extensions for the Ising model. First we show the exact evolution of the system for each of the quantum extensions. This is achieved by considering a small chain length of $N=10$ spins. Defining as initial ensemble, $\varrho_{0}$, the statistical mixture containing only configurations with zero magnetization, the time evolution of the chain is obtained by the successive application of the uniform mixture of map $\Lambda_{\mathcal{Q}}^{(q)}$, Eq.(14), applied to all the sites along the chain.
As we will see below the different quantum extensions modify the behavior of certain observables. What is more important for us is to analyze the possibility that quantum effects can accelerate and decelerate the process to reach the equilibrium subspace. An effect that has been reported previously [4].
Coherence.
To better understand the mechanism by which the relaxation process is modified we first study the coherence of the system, which is calculated in terms of the L1-norm
$\mathcal{C}(t)$ $=\sum_{i\neq j}\left|\varrho_{ij}(t)\right|\,,$ (18)
where $\varrho(t)$ is the density matrix that describes the state of the chain at time $t$. Recently it has been shown that coherence can be understood as a quantum resource that can be used to perform quantum tasks [5], in analogy to how the entanglement is the necessary resource to implement quantum protocols, as the teleportation, coherence is the necessary resource to modify the behavior of the relaxation process.
In figure (1) we observe the dynamics of the coherence as a function of time, in units of MC steps, for a chain of length $N=10$ and for the different quantum extensions $\{H_{0},H_{1},H_{2},S_{2},S_{1},S_{0}\}$, defined in Eqs.(14-17). At the beginning of the evolution the local map, $\Lambda_{\mathcal{Q}}^{(q)}$ introduce coherence along the system by the action of the unitary operations, $H$ and $S$. Then, the relaxation process start to dominate the dynamics and we observe the decaying of the coherence, after a couple of MC steps, with different decay rates for the different quantum extensions.
Figure 1: Coherence $vs$ $t$ (in units of MC steps). The different curves corresponds to different quantum extensions proposed above, see Eqs.(14-17). All cases are for a chain of $N=10$ spins.
Figure 2: Purity $vs$ $t$ (in units of MC steps). The different curves corresponds to different quantum extensions proposed above, see Eqs.(14-17). All cases are for a chain of $N=10$ spins.
Figure 3: Domains $vs$ $t$ (in units of MC steps). The different curves corresponds to different quantum extensions proposed above, see Eqs.(14-17). All cases are for a chain of $N=10$ spins.
@@
Figure 4: Probability to reach the equilibrium $vs$ $t$ (in units of MC steps). The different curves corresponds to different quantum extensions proposed above, see Eqs.(14-17). All cases are for a chain of $N=10$ spins.
One is tempted to think that the more coherences in the system, the faster the relaxation process finds the state (subspace in our case) of equilibrium, as typically happens in quantum search protocols [17, 18, 19]. However, as we will see below this is not what happens in our model.
@@
it takes the maximum value of 1 when the system is in a pure state and the minimum value of $1/d$, when the system is at the maximum statistical mixture, where $d=2^{N}$ is the dimension of the Hilbert space. The maximum statistical mixture refers to a diagonal density matrix, where the system is found with the same probability, $1/d$, in any of the classical configurations. In classical dynamics, the relaxation process sends the system to one of two possible equilibrium configurations, $|\bm{0}\rangle$ when all spins point up or $|\bm{1}\rangle$ when all point down. In quantum dynamics, the relaxation process sends the system to a statistical mixture of these two configurations which we denote as, $\varrho(t\rightarrow\infty)=p{|\bm{0}\rangle\langle\bm{0}|}+(1-p){|\bm{1}\rangle\langle\bm{1}|}$ (for $p\approx 1/2$). So for long times the purity saturates at the value 1/2, while its initial value is close to the minimum value since we start with a statistical mixture of all configurations with zero magnetization. The dynamics of the purity is shown in Fig.2 where the different curves correspond to the different quantum extensions. The yellow curve with empty squares corresponds to the classical dynamics, where followed by the local map $\Lambda_{\mathcal{Q}}^{(q)}$, a projective measurement is applied that destroys all possible coherences created in the previous step. It is interesting to note that in the three quantum evolutions $\{H_{0},H_{1},H_{2}\}$ the purity grows slower than the rest of the quantum versions, even slower than in the classical evolution. Which is paradoxical, since these three quantum extensions are the ones that produce the most coherences along the evolution, as observed in Fig.1. On the other hand, in the three quantum extensions $\{S_{0},S_{1},S_{2}\}$ the purity grows faster than in the classical evolution, with $S_{0}$ being the fastest, followed by $S_{2}$ and finally $S_{1}$. This is consistent with the degree of coherence in each of these quantum evolutions, see Fig. 1.
Decay of domain walls.
Another observable that helps us to understand the relaxation process is the decay of the number of domain walls. In a spin system like ours, a domain wall is understood as the interface between two spins pointing in opposite directions. Since at the beginning of evolution we start in a disordered phase the number of domain walls is maximal. As the spins begin to cluster into larger domains pointing in the same direction, the number of walls decreases as, $t^{-1/z}$. The parameter $z$ is known as the domain growth exponent. In the quantum version, to analyze the decay of domain walls we define the following operator: $\hat{D}_{W}=\sum_{j}n_{D}(j){|j\rangle\langle j|}$, where $n_{D}(j)$ is the number of domain walls in the $j$th configuration. Thus, we can calculate the expectation value of this observable as
The behavior of this observable is described in Figure 3, for the different quantum evolutions together with the classical one. For short times all quantum versions coincide with the classical evolution. After a couple of MC steps, the quantum evolutions separate and again cluster $\{H_{0},H_{1},H_{2}\}$ remaining above the classical curve, indicating that in this case the domain growth exponent is larger. While the evolutions $\{S_{0},S_{1},S_{2}\}$ are below the classical curve, showing again an acceleration in the relaxation process.
Probability to reach equilibrium.
To conclude this section we analyze the probability that the system reaches the equilibrium subspace. Recall that in the classical version the system, in the limit $t\rightarrow\infty$, remains in one of the two steady states, $|\bm{0}\rangle$ or $|\bm{1}\rangle$. On the contrary in the quantum evolution the system finish in a statistical mixture of these two states. The probability to reach the equilibrium, $P_{\text{Eq}}(t)$ can be defined in terms of the density matrix $\varrho(t)$ and the projector to the equilibrium subspace $\hat{P}_{\text{Eq}}={|\bm{0}\rangle\langle\bm{0}|}+{|\bm{1}\rangle\langle\bm{1}|}$ as follows
Figure 4 shows the behavior of the equilibrium probability as a function of time (in MC units). At the beginning of the evolution all curves, classical and quantum, behave similarly. After a few MC steps we see how the quantum curves move away from the classical one. As we saw above, the three curves $H_{0},H_{1},H_{2}$ with more coherence along the evolution are the ones that slow down the relaxation process, even slower than the classical process. While the three quantum versions $S_{0},S_{1},S_{2}$ make the relaxation process more efficient.
According to $\mathcal{P}(t)$, $\langle\hat{D}_{W}\rangle(t)$ and $P_{\text{eq}}(t)$ the most efficient quantum extension is $S_{0}\equiv$ SYH and $H_{0}\equiv$ HAD the most inefficient to accelerate the relaxation process. However, by ordering the different quantum evolutions from the most coherent to the least coherent, $\mathcal{C}(H_{0})>\mathcal{C}(H_{2})>\mathcal{C}(H_{1})>\mathcal{C}(S_{0})>\mathcal{C}(S_{2})>\mathcal{C}(S_{1})$, suggesting that the L1-norm is not the best candidate to describe quantum effects on macroscopic observables.
IV.1 Coherence
IV.1.1 Exact evaluation of the density matrix
We consider the evolution of the system in terms of the density matrix $\varrho(n)$ in the space of configurations (for $N$ spins, $2^{N}$ configurations). To generate the following figures, we classify all elements $\varrho_{ij}(n)$ with respect to three Hamming distances: the distance $d_{\rm H}(i,q_{\rm d})$ of the config. $i$ to the all-spins-down state $i_{\rm d}$, the distance $d_{\rm H}(j,q_{\rm d})$, and the Hamming distance $d_{\rm H}(i,j)$ between the configurations $i$ and $j$. Every class of matrix elements is then characterized by the triple $[d_{\rm H}(i,q_{\rm d}),d_{\rm H}(j,q_{\rm d}),d_{\rm H}(i,j)]$. The figures 5 and 6 show that average value of the matrix elements belonging to these classes.
For our simulations, the initial state is the uniform mixture of all configurations with magnetization zero. The density matrix then only has diagonal elements, and since all states have magnetization zero, their distance to the all-spins-down state ist $N/2$. Hence, in the space of distances, these matrix elements correspond to the single point $(N/2,N/2,0)$.
After one elementary operation, from any of the zero-magnetization configurations, one can arrive at a new state flipping a single spin. If the operation is done with certainty, we “move” along the diagonal of the density matrix towards the all-spins up or the all-spins down state. In this case, the distances for $i$ and $j$ remain equal and either increase or decrease by one unit. Alternatively, a unitary operation is applied, which produces non-diagonal elements. Now the distance between $i$ and $j$ increase by one unit.
In Figs. 5 and 6, we compare the relaxation dynamics when using $S_{0}$-gates (upper row) and $H_{0}$-gates (lower row).
Figure 5: Coherences after one MCS (Monte Carlo time step). $S_{0}$ (top row); $H_{0}$ (bottom row).
In Fig. 5, we analyse the density matrix after one Monte Carlo time step – this means $N$ elemental operations. In this case, it is possible that the system reaches one of the two minimum-energy configurations. However, as one can see in the first column (diagonal elements), the system arrives at those states $d_{\rm H}(i,j_{0})=0,N$ with very low probability, only. Here, almost no difference can be observed between use of $S_{0}$-gates (upper row) or $H_{0}$-gates (lower row). For the non-diagonal elements, the situation is diffent. In all cases (column $2-4$) differences are observable, most clearly when $d_{\rm H}(i,j)=2,3$. Note that in the $S_{0}$-case, the pattern are strictly symmetric (even distances) or anti-symmetric (odd distances) with respect to the off-diagonal connecting the points $(0,N)$ (upper left corner) and $(N,0)$ (lower right corner). This symmetry is lost in the $H_{0}$-case. Note also the slightly different scales for the average values of the matrix non-diagonal elements which implies that the coherences are somewhat larger in the $H_{0}$ case.
Figure 6: Coherences after the respective half time. $S_{0}$ (top row); $H_{0}$ (bottom row).
Fig. 6 shows the same quantities but at a later time, $t_{1/2}$, when the probability to find the system in the minimum energy subspace is equal to one half. For the $H_{0}$ case this is at a much later time ( …) than for the $S_{0}$ case. Again in the distribution for the diagonal elements it is hard to observe any differences. But for the non-diagonal elements, the amount of coherence is much larger for the $H_{0}$ case than the $S_{0}$ case (almost twice as much). We also observe that the symmetry in the pattern for the non-diagonal elements remains (is absent) in the $S_{0}$ ($H_{0}$) case.
IV.1.2 Scaling
@@
In this subsection we consider the scaling behaviour for the measure of coherence, defined in Eq. (18) of the quantum extensions $S_{0}$ and HAD for both the short and long times. To study the finite size scaling we have simulated the chain of Ising spin system for a larger number of spins that includes spin chanins of $N=12,~{}14,~{}16,~{}18$ and 20 spins. Here we use an unraveling method [20, 21], to avoid the evolution of huge density matrices. Considering a chain of length $N$ we define as the basic unit of time a Monte-Carlo step (MC), which consist of the application of the map, $\Lambda_{\mathcal{Q}}^{(q)}$, $N$-times. This is the minimum unit of time needed to observe relevant changes in the dynamics of the chain.
@@
with the following values for the parameters: $\alpha=2.0\pm 0.02$, $k_{1}\simeq 3.02$ for and $H_{0}$ and $\alpha=1.91\pm 0.014$, $k\simeq 5.23$ for $S_{0}$.
@@
Let us determine the crossover time $t_{c}(N)$. This is the time at which the decay of coherence cross from the initial behaviour given by euqation 22 to the long-time behavior given by equation 23. At the crossover time $t_{c}$, one can write
@@
$t_{c}(N\rightarrow\infty)\sim(\lambda-\alpha)\frac{N}{k}log(N)$ (26)
Please note that the crossover time is a function of $N$, even for the large systems sizes, which indicates that behavior of coherence described by equation 22 will prevail for $N\rightarrow\infty$.
@@ V Conclusions
We found several quantum extensions for the classical relaxation process in the Ising spin chain by applying symmetry operations. They show properties intermediate between the $S_{0}$ and $H_{0}$ cases studied earlier.
Since the coherence is the crucial ingredient, responsible for the differences in the behavior of macroscopic observables, we investigate in detail its scaling behavior. We found different scaling behaviors for short and long-time scaling, where the crossover time is a function of system size, even for large $N$.

### v2

@@ IV Numerical simulations
In the previous section we introduce a method that allows to generate several quantum extensions. Here we analyzed the exact evolution of the Ising chain under the influence of those quantum effects.
@@
Considering a small chain length of $N=12$ spins and defining as initial ensemble, $\varrho_{0}$, the statistical mixture that containing only configurations with zero magnetization, the time evolution of the chain is obtained by the successive application of the uniform mixture of map $\Lambda_{\mathcal{Q}}^{(q)}$, Eq. (14), applied to all the sites along the chain as is described in Eq. 12. As we will see below, the different quantum extensions modify the behavior of certain observables. What is more important for us is to analyze the possibility that quantum effects can accelerate or decelerate the relaxation process to the equilibrium. A phenomenon that has been previously reported [8].
Probability to reach equilibrium.
In the classical dynamics of the Ising chain, in the limit $t\rightarrow\infty$, the system evolves to one of the two steady states, $|\bm{0}\rangle\equiv|00\ldots 0\rangle$ or $|\bm{1}\rangle\equiv|11\ldots 1\rangle$. In contrast in the quantum evolution the system finishes in a statistical mixture of these two states. The probability to reach the equilibrium, $P_{\text{Eq}}(t)$ can be defined in terms of the density matrix $\varrho(t)$ and the projector to the equilibrium subspace $\hat{P}_{\text{Eq}}={|\bm{0}\rangle\langle\bm{0}|}+{|\bm{1}\rangle\langle\bm{1}|}$ as follows
where, $\mbox{tr}[A]$, stands for the trace of the operator $A$. Figure 1-a shows the behavior of the equilibrium probability as a function of time (in MC units). At the beginning of the evolution all curves, classical and quantum, behave similarly. After a few MC steps we see how the quantum curves move away from the classical one. Interestingly the quantum curves separate into 2 groups defined by the columns in Eq. III which differ only by the operation $H$ or $\sigma_{x}H$ to the left of the tensor product symbol. In the first group which we identify with solid lines in Figure 1-a, it is clearly observed that the continuous creation and annihilation of coherences slows dawn the relation process towards equilibrium. However, a clear acceleration of the relaxation process is seen in the dashed curves. The extreme cases of deceleration and acceleration are defined by the quantum operations defined in Eq. 17 and Eq. 18 respectively.
Decay of domain walls.
Another observable that is widely used in classical studies of the relaxation process in spin systems is the decay of the number of domain walls. A domain wall is understood as the interface between two spins pointing in opposite directions. Since at the beginning of evolution we start in a disordered phase, the number of domain walls is maximal.
@@
Figure 1: Observables measured for a linear chain of 12 spins. Panel (a), Probability to reach the equilibrium $vs\,t$ computed using Eq. 20. Panel (b), Purity $vs\,t$ computed from Eq. 22. Panel (c), Average of domain walls $vs\,t$ see Eq. 21. Panel (d), Coherence $vs\,t$ see Eq. 23. In all cases the time is measure in units of MC steps. The different curves corresponds to the different quantum extensions obtained in III.
As the spins begin to cluster into larger domains pointing in the same direction, the number of walls decreases as, $t^{-1/z}$. The parameter $z$ is known as the domain growth exponent. In the quantum version, to analyze the decay of domain walls we define the following operator: $\hat{D}_{W}=\sum_{j}n_{D}(j){|j\rangle\langle j|}$, where $n_{D}(j)$ is the number of domain walls in the $j$th configuration. Thus, we can calculate the expectation value of this observable as
The behavior of this observable is described in Figure 1-b, for the different quantum evolutions together with the classical one. For short times, quantum versions coincide with the classical evolution, but for long times, we observe how the quantum versions depart from the classical one in groups, such as in the previous case.
@@
it takes the maximum value of 1 when the system is in a pure state and the minimum value of $1/d$, when the system is at the maximum statistical mixture, where $d=2^{N}$ is the dimension of the Hilbert space. The maximum statistical mixture refers to a diagonal density matrix, where the system is found with the same probability $1/d$ in any of the classical configurations. In classical dynamics, the relaxation process sends the system to one of two possible equilibrium configurations, $|\bm{0}\rangle$ when all spins point up or $|\bm{1}\rangle$ when all spins point down. In quantum dynamics, the relaxation process sends the system to a statistical mixture of these two configurations which we denote as $\varrho(t\rightarrow\infty)=p{|\bm{0}\rangle\langle\bm{0}|}+(1-p){|\bm{1}\rangle\langle\bm{1}|}$ (for $p\approx 1/2$). So for long times the purity saturates at the value 1/2, while its initial value is close to the minimum value since we start with a statistical mixture of all configurations with zero magnetization. The purity dynamics is shown in Figure 1-c where the different curves correspond to the different quantum extensions. The yellow dotted curve corresponds to the classical dynamics, where followed by the local map $\Lambda_{\mathcal{Q}}^{(q)}$, a projective measurement is applied that destroys all possible coherence created in the previous step. The purity behavior is consistent with the two previous cases. The same group of quantum extensions show an improvement in the relaxation process, while in the second group the relaxation process is even slower than in the classical case.
Coherence.
To better understand the mechanism by which the relaxation process is modified we study the coherence of the system, which is calculated in terms of the L1-norm
$\mathcal{C}(t)$ $=\sum_{i\neq j}\left|\varrho_{ij}(t)\right|\,,$ (23)
where $\varrho(t)$ is the density matrix that describes the state of the chain at time $t$. Recently, it has been shown that coherence can be understood as a quantum resource that can be used to perform quantum tasks [9], in analogy to how entanglement is the necessary resource to implement quantum protocols, such as teleportation, coherence is the necessary resource to modify the behavior of the relaxation process.
In figure 1-d we observe the dynamics of the coherence as a function of time, in units of MC steps. At the beginning of the evolution, the local map $\Lambda_{\mathcal{Q}}^{(q)}$ introduces coherence along the system through the action of operations defined in III. Then, the relaxation process starts to dominate the dynamics, and we observe the decaying of the coherence after a couple of MC steps, with different decay rates for the different quantum extensions.
One is tempted to think that the more coherence in the system, the faster the relaxation process finds the state (subspace in our case) of equilibrium, as typically happens in quantum search protocols [25, 26, 27]. However, as we have seen so far, this is not always true.
Scaling for the coherence measure
@@
We consider the scaling behavior for the measure of coherence, defined in Eq. (23) of the quantum extensions $S_{0}$ and $H_{0}$ for both short and long times. To study the finite size scaling we have simulated the Ising model for spin chains of length $N=12,~{}14,~{}16,~{}18$ and 20. Here we use an unraveling method [7, 28], to avoid the evolution of huge density matrices. Considering a chain of length $N$ we define as the basic unit of time a Monte-Carlo step (MC), which consist of the application of the map, $\Lambda_{\mathcal{Q}}^{(q)}$, $N$-times. This is the minimum unit of time needed to observe relevant changes in the dynamics of the chain.
@@
with the following values for the parameters: $\alpha=2.0\pm 0.02$, $k_{1}\simeq 3.02$ for $H_{0}$ and $\alpha=1.91\pm 0.014$, $k_{1}\simeq 5.23$ for $S_{0}$.
@@
Let us determine the crossover time $t_{c}(N)$. This is the time at which the decay of coherence cross from the initial behavior given by equation 24 to the long-time behavior given by equation 25. At the crossover time $t_{c}$, one can write
@@
$t_{c}\sim\frac{\lambda-\alpha}{k}\,N\,log(N)$ (28)
Note that the crossover time is a function of $N$, even for the large systems sizes, which indicates that behavior of coherence described by equation 24 will prevail for $N\rightarrow\infty$.
@@ V Conclusions
In this work, we constructed several different quantum extensions, derived from a real discrete subgroup of $U(2)$. These extensions include the two basic ones, $S_{0}$ and $H_{0}$, already discussed in Ref. [8].
We analyze the exact relaxation dynamics of the Ising model for $N=12$ spins, where we find that the behavior of characteristic macroscopic observables is different for the different quantum extensions. In particular, we study the time dependence of (i) the probability to find the system in an equilibrium state, (ii) the number of domain walls, (iii) the purity, and (iv) the coherence. In all cases, except for case (iv), the results are very similar within two antagonistic groups of quantum extensions. In the first group superpositions are generated with the gate $H$, in the second group with the gate $\sigma_{x}H$. By consequence the original extension $H_{0}$ belongs to the first group, where the relaxation process is slower than in the classical case and $S_{0}$ to the second, where it is faster.
We include two quantum measures, the purity and the coherence (quantified as proposed in Ref. [9]). We expect the coherence to be particularly relevant for a quantum extension to show a different behavior than the original classical process. This is because it measures the distance between the density matrix with superpositions and the one where all superpositions (coherences) are removed. It is therefore natural to expect that the deviation from the classical behavior is directly related to the amount of coherence in the system. This expectation is confirmed only partially, as the behavior is rather heterogeneous among the different quantum extensions we have studied.
In Appendix A, we analyze the evolution of the density matrix in the space of three collective coordinates. In this way, we hope to approach a possible explanation of the different behaviors of the quantum extensions. In fact, we found certain characteristic differences, in particular an unexpected symmetry in the $S_{0}$ case, absent for $H_{0}$. However, we are still far from understanding this phenomenon.
We then study the scaling behavior of the coherence for the two antagonist quantum extensions. We found that the finite size scaling form exists for both the short and long time. We not only observed different scaling behaviors for the short and the long time but also found the values of the critical exponent to be different in these two time regime. However, the values of the exponents are similar for both the quantum extensions $S_{0}$ and $H_{0}$. We calculated the crossover time which is a function of system size, even for large $N$.
@@ Acknowledgements.
Appendix A Exact evaluation of density matrix with respect to Hamming distances
We consider the evolution of the system in terms of the density matrix $\varrho(n)$ in the space of configurations (for $N$ spins, $2^{N}$ configurations). To generate the following figures, we classify all elements $\varrho_{ij}(n)$ with respect to three Hamming distances: the distance $d_{\rm H}(i,q_{\rm d})$ of the config. $i$ to the all-spins-down state $i_{\rm d}$, the distance $d_{\rm H}(j,q_{\rm d})$, and the Hamming distance $d_{\rm H}(i,j)$ between the configurations $i$ and $j$. Every class of matrix elements is then characterized by the triple $[d_{\rm H}(i,q_{\rm d}),d_{\rm H}(j,q_{\rm d}),d_{\rm H}(i,j)]$. The figures 4 and 5 show that the average value of the matrix elements belonging to these classes.
For our simulations, the initial state is the uniform mixture of all configurations with magnetization zero. The density matrix then only has diagonal elements, and since all states have magnetization zero, their distance to the all-spins-down state is $N/2$. Hence, in the space of the distances, these matrix elements correspond to the single point $(N/2,N/2,0)$.
After one elementary operation, from any of the zero-magnetization configurations, one can arrive at a new state flipping a single spin. If the operation is done with certainty, we “move” along the diagonal of the density matrix towards the all-spins up or the all-spins down state. In this case, the distances for $i$ and $j$ remain equal and either increase or decrease by one unit. Alternatively, a unitary operation is applied, which produces non-diagonal elements. Now the distance between $i$ and $j$ increases by one unit.
In Figs. 4 and 5, we compare the relaxation dynamics when using $S_{0}$-gates (upper row) and $H_{0}$-gates (lower row) for $N=10$ spins.
Figure 4: Coherences after one MCS (Monte Carlo time step) for $N=10$ spins. $S_{0}$ (top row); $H_{0}$ (bottom row).
In Fig. 4, we analyze the density matrix after one Monte Carlo time step – this means $N$ elemental operations. In this case, it is possible that the system reaches one of the two minimum-energy configurations. However, as one can see in the first column (diagonal elements), the system arrives at those states $d_{\rm H}(i,j_{0})=0,N$ with very low probability, only. Here, almost no difference can be observed between use of $S_{0}$-gates (upper row) or $H_{0}$-gates (lower row). For the non-diagonal elements, the situation is different. In all cases (column $2-4$) differences are observable, most clearly when $d_{\rm H}(i,j)=2,3$. Note that in the $S_{0}$-case, the pattern are strictly symmetric (even distances) or anti-symmetric (odd distances) with respect to the off-diagonal connecting the points $(0,N)$ (upper left corner) and $(N,0)$ (lower right corner). This symmetry is lost in the $H_{0}$-case. Note also the slightly different scales for the average values of the matrix non-diagonal elements which implies that the coherences are somewhat larger in the $H_{0}$ case.
Figure 5: Coherences after the respective half time for $N=10$ spins. $S_{0}$ (top row); $H_{0}$ (bottom row).
Fig. 5 shows the same quantities but at a later time, $t_{1/2}$, when the probability to find the system in the minimum energy subspace is equal to one half. For the $H_{0}$ case this is at a much later time than for the $S_{0}$ case. Again in the distribution for the diagonal elements it is hard to observe any differences. But for the non-diagonal elements, the amount of coherence is much larger for the $H_{0}$ case than the $S_{0}$ case (almost twice as much). We also observe that the symmetry in the pattern for the non-diagonal elements remains (is absent) in the $S_{0}$ ($H_{0}$) case.